# Ringing Thick Braneworld with Finite Extra Dimension

### v1

@@ I Introduction
The detection of gravitational waves in 2015 by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo LIGOScientific:2016aoc sparked renewed interest in black hole physics, particularly in the realm of quasinormal modes (QNMs) Berti:2009kk ; Cardoso:2016rao ; Kokkotas:1999bd ; Konoplya:2011qq ; Jusufi:2020odz ; Nollert:1999ji ; Cheung:2021bol . In general, QNMs refer to the characteristic modes of a dissipative system. It is well known that QNMs can reveal the characteristic information of a black hole, such as mass, charge, angular momentum, and so on. Additionally, QNMs play a crucial role in other fields of physics. For example, as mentioned in Ref. Berti:2009kk , QNMs have become a valuable tool for detecting extra-dimensional signals. There are two main approaches: investigating the QNMs of black holes on the branes Chakraborty:2017qve ; Prasobh:2014zea ; Dey:2020pth ; Hashemi:2019jlt ; Chen:2016qii ; Konoplya:2003dd ; Cardoso:2003vt ; Seahra:2004fg ; Seahra:2006tm ; Chung:2015mna ; Dey:2020lhq ; Banerjee:2021aln ; Mishra:2021waw ; Lin:2022hus , or directly researching the QNMs of brane models Seahra:2005wk ; Seahra:2005iq ; Tan:2022vfe ; Tan:2023cra .
@@
As early as 2005, Seahra found that in the RS-II model, for a bulk observer, there are metastable gravitational wave bound states (also called resonance states), while for an observer on the brane, these states can be viewed as a set of discrete QNM spectrum of decaying massive gravitons. Recently, some of us found that, similar to the thin brane case, a thick brane can also possess a series of separated QNMs Tan:2022vfe . Furthermore, we also discovered that those long-lived QNMs have a close relationship with the resonance modes of a thick brane Tan:2023cra . This paper aims to investigate the difference in QNMs between the thick brane with a finite extra dimension and that with an infinite extra dimension. We also discuss the stability of the brane under various types of perturbations.
@@ V Conclusion
Similar to the results obtained in other braneworld models, the QNMs of the thick brane are short-lived. Thus, these modes are unlikely to be detected in current ground-based gravitational wave detection experiments. We can only hope they will be discovered in stochastic gravitational wave detection experiments Koyama:2004cf . Due to the unique properties of extra dimensions, this study may provide a new way to detect the existence of extra dimensions.
@@ Acknowledgments
We would like to thank Yu-Peng Zhang, Chun-Chun Zhu and Wen-Yi Zhou for very useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12205129, No. 12147166, No. 12247101, and No. 12347111), the China Postdoctoral Science Foundation (Grant No. 2021M701529 and No. 2023M741148), the 111 Project (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University.

### v2

# Quasinormal ringing of thick braneworlds with a finite extra dimension

@@ I Introduction
The detection of gravitational waves in 2015 by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo LIGOScientific:2016aoc sparked renewed interest in black hole physics, particularly in the realm of quasinormal modes (QNMs) Berti:2009kk ; Cardoso:2016rao ; Kokkotas:1999bd ; Konoplya:2011qq ; Jusufi:2020odz ; Nollert:1999ji ; Cheung:2021bol . In general, QNMs refer to the characteristic modes of a dissipative system. It is well known that QNMs can reveal the characteristic information of a black hole, such as mass, charge, angular momentum, and so on. Additionally, QNMs play a crucial role in other fields of physics. For example, as mentioned in Ref. Berti:2009kk , QNMs have become a valuable tool for detecting extra-dimensional signals. There are two main approaches: investigating the QNMs of black holes on the branes Chakraborty:2017qve ; Prasobh:2014zea ; Dey:2020pth ; Hashemi:2019jlt ; Chen:2016qii ; Konoplya:2003dd ; Cardoso:2003vt ; Seahra:2004fg ; Kanti:2005xa ; Seahra:2006tm ; Kanti:2006ua ; Ishihara:2008re ; Chung:2015mna ; Bronnikov:2019sbx ; Dey:2020lhq ; Banerjee:2021aln ; Mishra:2021waw ; Lin:2022hus ; Bohra:2023vls ; Zinhailo:2024jzt , or directly researching the QNMs of brane models Seahra:2005wk ; Seahra:2005iq ; Tan:2022vfe ; Tan:2023cra .
@@
As early as 2005, Seahra found that in the RS-II model, for a bulk observer, there are metastable gravitational wave bound states (also called resonance states), while for an observer on the brane, these states can be viewed as a set of discrete QNM spectrum of decaying massive gravitons. Recently, some of us found that, similar to the thin brane case, a thick brane can also possess a series of separated QNMs Tan:2022vfe ; Tan:2024url . Furthermore, we also discovered that those long-lived QNMs have a close relationship with the resonance modes of a thick brane Tan:2023cra . This paper aims to investigate the difference in QNMs between the thick brane with a finite extra dimension and that with an infinite extra dimension. We also discuss the stability of the brane under various types of perturbations.
@@ IV.2 The case of $n=1$
To verify the analytical results (78), (79), and (80), we use the continued fraction method Leaver:1985ax ; Leaver:1990zz to calculate the QNMs of various perturbations and bulk fields for the case of $n=1$. Namely, we need to solve the Schrödinger-like equation (74) with the outgoing boundary conditions
$\psi(\bar{z})\propto\left\{\begin{aligned} e^{i\bar{m}\bar{z}}&,\quad\bar{z}\rightarrow\infty,\\ e^{-i\bar{m}\bar{z}}&,\quad\bar{z}\rightarrow-\infty,\end{aligned}\right.$ (81)
where $\bar{m}^{2}=\frac{m^{2}}{k^{2}}-\alpha$.
First, by performing a coordinate transformation $\tilde{z}=\tanh\bar{z}$, Eq. (74) and the corresponding boundary conditions (81) can be rewritten as
$\left\{(1-\tilde{z}^{2})\partial_{\tilde{z}}^{2}-2\tilde{z}\partial_{\tilde{z}}+\left[\frac{\bar{m}^{2}}{1-\tilde{z}^{2}}+\beta\right]\right\}\psi(\tilde{z})=0,$ (82)
and
$\psi(\hat{z})\propto\left\{\begin{aligned} &(1+\tilde{z})^{-i\bar{m}/2},&\tilde{z}\rightarrow-1,\\ &(1-\tilde{z})^{-i\bar{m}/2},&\tilde{z}\rightarrow 1,\end{aligned}\right.$ (83)
respectively. Then, following the continued fraction method proposed by Leaver Leaver:1985ax , the ansatz satisfying the boundary conditions (83) for the wave function $\psi(\tilde{z})$ is
$\psi(\tilde{z})=(1+\tilde{z})^{-i\bar{m}/2}(1-\tilde{z})^{-i\bar{m}/2}\sum_{l=0}^{\infty}a_{l}\tilde{z}^{l},$ (84)
where $a_{l}$ is the expansion coefficient. Substituting this into Eq. (82), a two-term recurrence relation can be obtained:
$\gamma_{l}a_{l+2}+\delta_{l}a_{l}=0,\quad l=0,1,2,\cdots$ (85)
where the recurrence coefficients $\gamma_{l}$ and $\delta_{l}$ are functions of $l$ and the parameter $\beta$:
$\gamma_{l}$ $=2+l-(2+l)^{2},$ (86)
$\delta_{l}$ $=(l^{2}-\beta+l(1-2i\bar{m})-\bar{m}(i+\bar{m})).$ (87)
In fact, the recurrence relation (85) can be written into the following matrix form:
$\begin{pmatrix}\delta_{0}&0&\gamma_{0}&&&\\ &\delta_{1}&0&\gamma_{1}&&\\ &&\delta_{2}&0&\gamma_{2}&\\ &&&\ddots&\ddots&\ddots\end{pmatrix}\begin{pmatrix}a_{0}\\ a_{1}\\ a_{2}\\ \vdots\end{pmatrix}=0.$ (88)
It can be seen that this matrix equation (88) is essentially an algebraic equation for the QNMs. Solving this linear system yields the QNMs, and this system has nontrivial solution if and only if the determinant of the coefficient matrix is zero.
Generally, solving the eigenvalue problem for this matrix equation involves dealing with an infinite-dimensional secular equation. To make the problem solvable, we can truncate it at a sufficiently large dimension, and the larger the dimension of the matrix, the more accurate the results. Fortunately, since this coefficient matrix is an upper triangular matrix, we can directly obtain:
$\delta_{0}=0\;,\;\delta_{1}=0\;,\;\delta_{2}=0\;,\;\cdots\;,\;\delta_{l}=0\;,\;\cdots$ (89)
namely,
$\bar{m}=\frac{1}{2}\left[-i(1+2l)\pm\sqrt{-1-4\beta}\right],\quad l=0,1,2,\cdots.$ (90)
Finally, through the relation $\bar{m}^{2}=\frac{m^{2}}{k^{2}}-\alpha$, we find that the results are completely consistent with those obtained from the analytical solution.
@@ V Conclusion
Similar to the results obtained in the RS-II thin brane, the QNMs of the thick brane are short-lived. It seems that current laser interferometer gravitational wave detectors have no way to detect these QNMs. But for the case of nonvanishing spatial three-momentum $p^{2}\neq 0$ or $n$ close to $1$, the QNMs will exhibit a slowly oscillating decay tail which can be seen in Fig. 8. According to Ref. Konoplya:2023fmh , these tails may have an impact on the stochastic gravitational wave detection experiments such as the Pulsar Timing Array experiment NANOGrav:2023gor ; Koyama:2004cf ; Caprini:2018mtu . Due to the unique properties of extra dimensions, this study may provide a new way to detect the existence of extra dimensions.
@@ Acknowledgments
We would like to thank Yu-Peng Zhang, Chun-Chun Zhu and Wen-Yi Zhou for very useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12475056, No. 12205129, No. 12247101, and No. 12347111), the China Postdoctoral Science Foundation (Grant No. 2023M741148), the 111 Project (Grant No. B20063), the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20240458), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University.